Time CF 0,5 0,6mln 1 6mlin 1,5 5mln 2 4mlrn a) Calculate a payback period for this investment: b) Calculate discounted payback period for this investment with 6% interest rate; c) Calculate NPV.
Answer #1
a) Time CF Cumulative CF
0 -11 -11
0.5 0.6 -10.4
1 6 -4.4
1.5 5 0.6
2 4 4.6
Payback period = 1+0.5*4.4/5 = 1.44
b) Time CF PVIF at 3% PV at 3% Cumulative CF
0 -11 1.00000 -11.00 -11
0.5 0.6 0.97087 0.58 -10.42
1 6 0.94260 5.66 -4.76
1.5 5 0.91514 4.58 -0.19
2 4 0.88849 3.55 3.37
3.37
Discounted payback = 1.5+0.5*0.19/3.55 = 1.53
c) NPV = 3.37 mln